OpenStudy (anonymous):

Given the domain {-3, 0, 6}, what is the range for the relation 2x + y = 4? {10, 4, -8} {-10, -4, 8} {2, 4, 10} {-2, 4, 16}

OpenStudy (anonymous):

Plugin each value of the domain into the function in place of 'x' to get the corresponding value of 'y'. This will give you the range

OpenStudy (anonymous):

Well yah but........

OpenStudy (anonymous):

For example if you plug in x = -3 in the equation 2x + y = 4, then you get 2(-3) + y = 4 -6 + y = 4 y = 4 + 6 y = 10 Hence 10 should be in the range. Can you do the same for the other values of x in the domain?

OpenStudy (anonymous):

K! thanx :)

OpenStudy (anonymous):

Did you get the answer it should be A

OpenStudy (anonymous):

Yah i put a thanks. It was right :)